OpenStudy (anonymous):

are these numbers complex? help please? -7, sqt16, sqt-16, sqt-15, 44, pi, 1.765, -1.5, sqt-6, 0,1

OpenStudy (anonymous):

all real numbers are also complex numbers

OpenStudy (radar):

\[\sqrt{-16},\sqrt{-15},\sqrt{-6}\]

OpenStudy (anonymous):

is e real?